OpenStudy (anonymous):

Find the vertex equals the absolute value of the quanity 3x+6 minus 4 of y is (2, -4) (-2, 4) (-2, -4) (-6, -4)

OpenStudy (helder_edwin):

\[ \large |3x+6|=y-4 \] ???

OpenStudy (anonymous):

y in front 3x+6 but everything else right

OpenStudy (helder_edwin):

\[ \large y=|3x+6|-4 \] ??

OpenStudy (anonymous):

yes

OpenStudy (helder_edwin):

it can be rewritten into \[ \large y+4=3|x+2| \] ok?

OpenStudy (anonymous):

ok well what would i do next to do simplify

OpenStudy (helder_edwin):

remember the equation of a parabola: \[ \large y-k=a(x-h)^2 \] has vertex \((h,k)\). it is the same here the vertex is in (-2,-4)

OpenStudy (anonymous):

thanks